The swinging pendulum has a period of 4.5 s, find the length of a swinging pendulum.


1
Expert's answer
2021-01-27T09:31:41-0500

By the definition of the period of the pendulum, we have:


"T=2\pi\sqrt{\dfrac{L}{g}},""L=\dfrac{T^2g}{4\pi^2}=\dfrac{(4.5\ s)^2\cdot 9.8\ \dfrac{m}{s^2}}{4\pi^2}=5.03\ m."